A retailer buys a batch of 50 items of a product for 10 000 €. He then sets a price so that the
selling price is 80 % higher than the buying price. When there is 40 % left of the batch, the
retailer lowers the price and manages to sell off the rest for a reduced selling price of
171 €/item.
a) Calculate the discount percentage relative to the initial selling price.
b) Calculate the profit percentage for the whole batch.
Give your answers with an accuracy of one decimal.
.
.
I TRIED IT THIS WAY
Let's resume what he did.
he bought 50 items for 10000€ means 1 item for 200€
when he set the selling price, it was 200*1.80 = 360€ for one item.
60% out of 50 items were sold by 360€/item...that gives us 360*30items = 10800€.
40% that is 20 items were sold by 171€*20 = 3420€.
a)
the discount percentage relative to the initial selling price: is the discount from 200€/item to 171€/item
200€-----------100%
171€-----------x%
x% = 85.5%
The discount is 100%-85.5%= 14.5%
b)
10800€+3420€ = 14220€ for the whole batch
that's:
10000€-------100%
14220€-------y%
y% = 142.2%
the profit percentage for the whole batch is 42.2% ?

Answer by MathLover1(20855)


you are correct............

For the profit percentage, the initial cost is 10,000€,
that's 200€ each , so an 80% increase would be 360€ each.
He sells 40%, or 30 items at 360€ = 10,800€
He sells 60%, or 20 items at 171€ = 3,420€ total = 14,220€
So the profit percent would be 14,220 - 10,000 = 4220/10,000=0.422
or 42.2%